1) 3^2x-4(3^x-2)+1=0
2) 125^x²-0.2(25^x)²=0
1
Expert's answer
2020-01-17T11:46:55-0500

1)

"3^{2x}-4(3^x-2)+1=0,"

substitution "t=3^x," "t^2=3^{2x},"

t2-4(t-2)+1=0

t2-4t+9=0

D=16-36= -20<0,

there is no real solution.


2)

"125^{x^2}-0.2(25^x)^2=0"

"125^{x^2}=5^{-1}25^{2x}"

"5^{3x^2}=5^{4x}5^{-1}"

"3x^2=4x-1"

3x2-4x+1=0

D=16-12=4

x=(4-2)/(2*3) and x=(4+2)/(2*3),

x=1/3 and x=1.